# Detailed nuclear structure calculations for coherent elastic neutrino-nucleus scattering

### v1

@@ Abstract
Any discovery of ‘new physics’ in the neutrino sector first requires a precise prediction of the expected Standard Model cross section. Currently, Coherent Elastic neutrino-Nucleus Scattering (CEvNS) experiments are statistics limited. However, as new and future experiments scale up, it will be necessary to improve the theoretical predictions. Here we review the calculation of the CEvNS cross section in a consistent theory of hadronic currents and compute the relevant nuclear form factors using the nuclear shell model. The uncertainty on the form factors is explored by repeating the calculation for various shell model interactions and with Skyme-Hartree-Fock evaluations of the Weak-charge radii. We then refine the Standard Model predictions for the recent experimental results of the COHERENT experiment. We find that our cross sections are in good agreement with previous predictions, but with significantly smaller uncertainties - by up to a factor of 10. Near-future CEvNS experiments will meaningfully benefit from improved predictions through an increased sensitivity to new-physics signals.
@@ I Introduction
The first detection of neutrinos via CEvNS was performed by the COHERENT experiment using a small (14.6 kg) cesium-iodide detector [1]. This measurement was quickly followed up by a detection using a liquid argon detector [2]. The next planned measurement will be with an array of germanium detectors (18 kg) [3]. All of these measurements were carried out at the Spallation Neutron Source (SNS) at Oak Ridge National Laboratory. There, neutrinos with energies up to $\sim$50 MeV are created from pions decaying at rest within the mercury spallation target. Together, these measurements have observed hundreds of CEvNS events and are broadly consistent with the Standard Model expectation for the process.
CEvNS has many potential applications, both practical and fundamental. While the practical applications, such as nuclear reactor monitoring [4], require further detector development before they can be realized, advances in nuclear [5, 6, 7] and fundamental physics are already being realized with current experiments [8, 9, 10, 11]. Notably, CEvNS experiments are the leading probe of some realizations of sterile neutrinos and non-standard neutrino interactions [12]. CEvNS could be detected in dark matter experiments and potentially provide the first warning of a supernova signal [13]. Key to maximising the effectiveness of these applications is having precise predictions for the scattering rate. That is, if one is interested in detecting deviations from the Standard Model at the sub-percent level, then a commensurate level of precision is required for the theoretical predictions.
@@
In this work we calculate the CEvNS cross sections for nuclear targets used by COHERENT, including the nuclear structure effects, which are typically ignored. To this end, we use large-scale nuclear shell model calculations to derive all relevant nuclear form factors. Using several different nuclear Hamiltonians, we evaluate the uncertainty in the cross section from the nuclear structure effects. These results are then applied to improve the theoretical predictions of the CEvNS event rates that COHERENT has observed to date.
@@ II CEvNS cross section
$E_{\nu}^{\rm{max}}=\frac{2E_{\nu}^{2}}{m_{T}+2E_{\nu}}.$ (2)
The cross section is typically calculated by treating the nucleus as composed of point-like nucleons and accounting for their spatial extent within the nucleus via a form factor. This results in a standard form of the CEvNS cross section
@@
where $Q_{\text{w}}=Q_{p}Z+Q_{n}(A-Z)$ is the weak nuclear charge in terms of the proton and neutron weak charges and $F(q)$ is the nuclear form factor, typically taken to be of the Helm [20] or Klein-Nystrand [21] form.
@@
×$\times\left(Q_{p,\nu_{i}}F_{p}(q^{2})+Q_{n}F_{n}(q^{2})\right)^{2}.$ (4)
Here the flavor dependent charges are taken from [23]:
$Q_{n}$ $=$ $-1.02352(25)$
$Q_{p,\nu_{e}}$ $=$ $0.0747(34),\,\,\,\,Q_{p,\nu_{\mu}}=0.0582(34).$ (5)
Using this equation requires knowledge of the proton and neutron distributions within the nucleus. While the former can be measured well due to its electric charge, the latter is not well constrained experimentally. Given the much larger weak charge of the neutron, the relatively unknown neutron form factor thus contributes the largest uncertainty to determinations of the CEvNS cross section.
III Nuclear form factors
@@
The density distribution used, i.e. electric charge-, weak charge- or mass-density, results in different form factors (and rms radii) which are appropriate for use in different scattering processes. Neutral-current neutrino scattering proceeds through the coupling of the Z-boson to the weak charge of the nucleus. Therefore CEvNS requires the use of weak-charge form factors which in turn require knowledge of the weak-charge density. We note that some past work has used experimentally determined electric-charge rms radii to construct form factors for CEvNS, but this can differ significantly from the weak-charge radii [26].
It is possible to measure the weak charge distribution through parity-violating electron scattering, which has been performed at Jefferson lab by the PREX and CREX experiments, on lead-208 and calcium-48 respectively. However, these experiments have only measured the weak charge at a single momentum transfer: $q=0.76$ fm-1 and 1.28 fm-1, respectively [27, 28]. Additionally, these experiments are difficult and unlikely to be performed on every isotope relevant to CEvNS. Therefore, it is necessary to use indirect knowledge of the weak-charge distribution from e.g. the electric charge distributions, or theoretical modelling.
In this work we will compare several methods for modelling the CEvNS form factors and determine their effect on the resulting cross sections. This includes theoretical modelling using both Skyrme-Hartree-Fock (SHF) calculations and the nuclear shell model. The SHF results will allow us to accurately model the nucleon distributions, from which rms radii are computed. The shell model calculations will allow us to include sub-leading nuclear effects through the chiral EFT formalism of [15]. For comparison with previous work in the literature, we will also perform a data-driven calculation based on electric-charge radii.
When modelling the nucleus as a bound state of point-like nucleons, the neutron and proton point radius distributions $R_{n},p$ are obtained from
@@
$R_{p}^{2}=\frac{1}{Z}\int d^{3}r\,r^{2}\rho_{p}(\mathbf{r}),$ (9)
respectively. Given that stable nuclei exhibit a common, almost uniform, saturating density it is possible to parameterise their density distributions (to very good accuracy) with fits to distributions of just two parameters: one which defines the size of the nucleus and another which defines the thickness of the ‘skin’ region at the surface. At low momentum transfers, the form factor remains relatively insensitive to differences in these parameterizations. To fit observation to a density distribution, it can be sufficient to equate a measured or theoretically determined rms radius with that of the parameterized distribution. We will make use of this matching process below.
III.1 From electric-charge densities
@@
The most reliable access we have to the nuclear size and shape is via electromagnetic probes which couple to the electric charge distribution. Protons, being charged, contribute more to this distribution than the neutrons do. However, the reverse is true for the weak-charge distribution where neutrons dominate in their contribution. This makes the electric charge distribution an imperfect surrogate for the weak-charge distribution. To account for this difference, previous works have made some assumptions about the neutron distribution based on the proton distribution. In [22] the neutron rms radius was chosen to lie between the proton rms charge radius and an additional 0.3 fm. In [29], a fit is made to the proton-neutron radii difference data, however the large scatter in the data used in the fit tends to exaggerate the uncertainty for any one nucleus.
@@
In this work we will follow the method of [22]. For our electric-charge inspired distributions we will use the Klein-Nystrand (KN) [21] model which has been previously used in CEvNS research (including by the COHERENT collaboration). The Klein-Nystrand form factor is based on both a Yukawa potential of range $a_{k}$ and a hard sphere distribution with radius $r_{A}$. It has the form
$F_{\text{KN}}(q)=3\frac{\sin(qr_{A})-qr_{A}\cos(qr_{A})}{(qr_{A})^{3}}\left[1+(qa_{k})^{2}\right]^{-1},$ (10)
where $a_{k}=0.7$ fm and $r_{A}$ is the nuclear radius. The nuclear radius is obtained by equating the rms radius radius for the distribution:
$\langle r^{2}\rangle_{KN}=\frac{3}{5}r_{A}^{2}+6a_{k}^{2}.$ (11)
@@
with the proton rms charge radius. The proton rms charge radii for the target nuclei $\sqrt{\langle r_{p}^{2}\rangle}\equiv r_{\text{rms}}^{p}$ can be obtained from various experimental sources. The values we adopt are shown in Tab. 1. The neutron rms radii are then taken to lie in the range $r_{\text{rms}}^{p}\leq r_{\text{rms}}^{n}\leq r_{\text{rms}}^{p}+0.3$ fm for Ge, I and Cs, whilst for Ar the upper limit is taken to be $r_{\text{rms}}^{p}+0.2$ fm. We use this range to bracket the uncertainty in the neutron distribution.
Table 1: Experimental charge rms radii $r_{\text{rms}}^{p}$ values (in fm) for 40Ar, 70,72,73,74,76Ge, 127I and 133Cs, obtained from [30].
40Ar Ge 127I 133Cs
70 72 73 74 76
III.2 From Skyrme-Hartree-Fock distributions
@@
To model the weak-charge distribution of the nucleus more accurately, we perform Skyrme-Hartree-Fock calculations to find the point-like nucleon distributions within the nucleus. While we have experimental probes of the electric-charge distributions, which dominantly probes the protons, for consistency we will also use the theoretical values for $R_{p}$ in this calculation. The weak radius of the nucleus can then be calculated from:
$R_{W}^{2}=\frac{1}{Q_{\text{w}}}\left(ZQ_{p}(R_{p}^{2}+r_{\text{w},p}^{2})+(A-Z)Q_{n}(R_{n}^{2}+r_{\text{w},n}^{2})\right)$ (12)
where $r_{\text{w},p}=1.545(17)$ fm and $r_{\text{w},n}=0.8434(12)$ fm are the weak radii of the proton and neutron, respectively.
@@
The nucleon densities were evaluated with the skyax solver [31] assuming axial symmetry for several parametrizations of the Skyrme functional, namely SkM, SLy6, SV-min and UNEDF1 as defined in skyax. Pairing correlations are treated at the BCS level with a “density-dependent delta interaction” (DDDI) for SV-min and SkM, and with a “volume delta interaction” (VDI) for SLy6 and UNEDF1. The final theoretical values of $R_{p,n}$ are obtained from the average of the results for each interaction, while the standard deviations are used to evaluate the uncertainties. The full results for each interaction are compiled in the appendix (Table 6). From this we find the average and standard deviation for each isotope and use it to compute weak radii from Eq. (12). The results are summarised in Table 2.
@@
For simplicity we then parameterize the distribution via the symmetrized 2-parameter Fermi (Fermi) model [32], which has been shown to provide a reasonable fit to similar calculations [26]. The Fermi distribution is defined by:
@@
The (normalised) form factor is then:
@@
×$\times\frac{\pi qa}{\sinh(\pi qa)}\frac{1}{1+(\pi a/c)^{2}}$
@@
Table 2: Skyrme-Hartree-Fock predictions of rms point-like proton and neutron radii and the resulting predicted RMS Weak radius.
@@
IV Detailed nuclear treatment
@@
The full, bottom-up, calculation starts with the hadronic currents within the nucleus and using effective field theory, derives the Wilson coefficients for the nuclear couplings. This task has been performed for the CEvNS process, in full generality, in [15]. The resulting cross section is much like the traditional form of Eq.(3), with an additional term for the axial form factor:
@@
However, the full form of the nuclear form factors go beyond the density distributions discussed previously. The Weak form factor takes the form
@@
where $m_{N}$ is the nucleon mass and the remaining constants are provided in [15]. The weak form factor hence depends on the coherent and semi-coherent responses $F^{M}_{N}$ and $F^{\Phi^{\prime\prime}}_{N}$, respectively, where $N=\{p,n\}$ denotes protons and neutrons. The axial-vector form factor has the form
@@
The correction to the leading SD coupling for the transverse SD response $\delta^{\prime}(\mathbf{q}^{2})$ is given in [15]. In this more-detailed treatment of the nucleus, the full Weak radius receives additional contributions from spin-orbit effects and the strange electric-charge distribution. It takes the form
@@
To apply this formalism and calculate cross sections for a given target we will employ the nuclear shell model to compute the nuclear responses above.
IV.1 Evaluating the nuclear responses
Figure 2: Spin-independent weak form factors for CEvNS in the four targets relevant to current experiments. Bands show the error due to uncertainties in the neutron distribution (Charge) and from using different interactions (SHF and Shell model).
To compute the weak and axial-vector nuclear form factors in Eqs. (18) and (19), we need to evaluate the $\mathcal{F}^{M}_{N}$, $\mathcal{F}^{\Phi^{\prime\prime}}_{N}$ and $\mathcal{F}^{\Sigma^{\prime}}_{N}$ nuclear responses. The nuclear responses originate from a multipole decomposition of the lepton-nucleus amplitude, developed in [33, 34]. Here, they are defined as:

### v2

@@ Abstract
Any discovery of ‘new physics’ in the neutrino sector will first require a precise baseline prediction of the expected Standard Model cross section. While Coherent Elastic neutrino-Nucleus Scattering (CEvNS) experiments are currently statistics limited, upcoming and future experiments at larger scales will be systematics limited. At this point it will be necessary to improve the theoretical predictions to improve the experimental sensitivity. Here we review the calculation of the CEvNS cross section in a consistent theory of hadronic currents and compute the relevant nuclear form factors using the nuclear shell model. The uncertainty on the form factors is explored by repeating the calculation for various shell model interactions and with Skyme-Hartree-Fock evaluations of the Weak-charge radii. We then refine the Standard Model predictions for the recent experimental results of the COHERENT experiment. We find that our cross sections are in good agreement with previous predictions, but with significantly smaller uncertainties. Near-future CEvNS experiments will meaningfully benefit from improved predictions through an increased sensitivity to new-physics signals.
@@ I Introduction
CEvNS has many potential applications, both practical and fundamental. While the practical applications, such as nuclear reactor monitoring [4], require further detector development before they can be realized, advances in nuclear [5, 6, 7] and fundamental physics are already being realized with current experiments [8, 9, 10, 11]. Notably, CEvNS experiments are the leading probe of some realizations of sterile neutrinos and non-standard neutrino interactions [12]. CEvNS will also be detected in dark matter experiments, where it is an irreducible background to dark matter searches, giving rise to the so-called neutrino floor [13] or fog [14]. Recently, XENONnT and PandaX reported the first indication of CEvNS due to 8B solar neutrinos [15, 16]. Key to maximizing the effectiveness of these applications is having precise predictions for the scattering rate. That is, if one is interested in detecting deviations from the Standard Model at the sub-percent level, then a commensurate level of precision is required for the theoretical predictions.
The dominant systematic uncertainty in CEvNS rate predictions comes from the neutrino flux, which is presently $\sim 10\%$ for the pion-decay-at-rest neutrinos produced at the SNS. However, future dedicated measurements by COHERENT aim to reduce this significantly, to around 2-3% [17]. With this improvement, cross section uncertainties, currently around 2% [2], will provide a significant contribution to the overall systemic uncertainty in the predicted event rates. Relatively little attention has been given to the theoretical uncertainties of the cross section. With CEvNS being a low-energy process, it is expected that predictions that treat the nucleus as an elementary particle are sufficient. The nuclear structure is then only accounted for through a simple charge form factor. Recently, the cross section calculation has been revisited and a modern accounting of nucleon currents applied [18]; however, this formalism has not yet been adopted for use in predicting experimental observables. That said, cross section calculations using alternative approaches, using coupled cluster theory [19] (for argon only) and Skyrme-Hartree-Fock theory [20] (for carbon, oxygen, argon, iron and lead) have been performed.
In this work we present theoretical predictions for the CEvNS cross sections, based on detailed nuclear structure calculations, for targets used by the COHERENT experiment (argon, germanium, cesium and iodine). In section II we review the standard calculation of the CEvNS cross section and present two methods to derive the nuclear form factors. In section III we revise the full CEvNS cross section calculation (following reference [18]) and present our computations of the relevant nuclear form factors based on the nuclear shell model. In section IV we use our form factors to improve the theoretical predictions of the CEvNS flux averaged cross sections and event rates for the COHERENT experiment. Finally, in section V we conclude.
@@ II CEvNS cross section
$E_{R}^{\rm{max}}=\frac{2E_{\nu}^{2}}{m_{T}+2E_{\nu}}.$ (2)
The cross section is typically calculated by treating the nucleus as composed of point-like nucleons and accounting for their spatial extent within the nucleus via a form factor. This results in a standard expression for the CEvNS cross section
@@
where $Q_{\text{w}}=Q_{\text{w}}^{p}Z+Q_{\text{w}}^{n}(A-Z)$ is the weak nuclear charge in terms of the proton and neutron weak charges and $F(q)$ is the nuclear form factor, typically taken from the Helm [22] or Klein-Nystrand [23] model.
@@
×$\times\left(Q_{\text{w}}^{p,\nu_{i}}F_{p}(q^{2})+Q_{\text{w}}^{n}F_{n}(q^{2})\right)^{2}.$ (4)
Here we take the flavor dependent charges from [25]:
$Q_{\text{w}}^{n}$ $=$ $-1.02352(25)$
$Q_{\text{w}}^{p,\nu_{e}}$ $=$ $0.0747(34),\,\,\,\,Q_{\text{w}}^{p,\nu_{\mu}}=0.0582(34),$ (5)
and use them in all further calculations. Using Eq. (II) requires knowledge of both the proton and neutron form factors. While the former can be inferred from the electric charge distribution, the latter is not well constrained experimentally. Given that the weak charge of the neutron is much larger than the proton’s, we can deduce that the dominant contribution to the uncertainty of the CEvNS cross section comes from the neutron form factor.
Table 1: Experimental charge rms radii $r_{\text{rms}}^{p}$ values (in fm) for 40Ar, 70,72,73,74,76Ge, 127I and 133Cs, from [28].
40Ar Ge 127I 133Cs
70 72 73 74 76
II.1 Nuclear form factors
@@
The density distribution used, i.e. electric charge-, weak charge- or mass-density, results in different form factors (and rms radii) which are appropriate for use in different scattering processes. When modelling the nucleus as a bound state of point-like nucleons, the neutron and proton point radii, $R_{n,p}$, are obtained from their respective distributions via
@@
and
$R_{p}^{2}=\frac{1}{Z}\int d^{3}r\,r^{2}\rho_{p}(\mathbf{r}).$ (9)
Neutral-current neutrino scattering proceeds through the coupling of the Z-boson to the weak charge of the nucleus. Therefore CEvNS requires the use of weak-charge form factors which in turn require knowledge of the weak-charge density. It is possible to measure the weak charge distribution directly through parity-violating electron scattering. This has been performed at Jefferson lab by the PREX and CREX experiments, on 208Pb and 48Ca targets, respectively. However, these experiments have only measured the weak charge at a single momentum transfer: $q=0.76$ fm-1 and 1.28 fm-1, respectively [29, 30]. Additionally, these experiments are difficult and unlikely to be performed on every isotope relevant to CEvNS. Therefore, it is necessary to use indirect knowledge of the weak-charge distribution from e.g. the electric charge distributions, or theoretical modelling. While electric-charge rms radii are measured to high-precision, they can differ significantly from the weak-charge radii [31] and form factors [20]. Therefore they cannot be used to produce accurate CEvNS cross sections.
In this work we will compare several methods for modelling the CEvNS form factors and determine their effect on the resulting cross sections. This includes theoretical modelling using both Skyrme-Hartree-Fock (SHF) calculations and the nuclear shell model. The SHF results will allow us to model the nucleon distributions accurately, from which rms radii are computed. The shell model calculations will allow us to include sub-leading nuclear effects through the chiral EFT formalism of [18]. For comparison with previous work in the literature, we will also perform a data-driven calculation based on electric-charge radii.
II.2 Empirically derived form factor
@@
The most reliable measurements of nuclear size come from probes of the nuclear electric-charge distribution, e.g., from electron scattering data. Protons, being charged, contribute more to the distribution than the neutrons do. However, the reverse is true for the weak-charge distribution (where the neutron contribution dominates), making the electric-charge distribution an imperfect surrogate for the weak-charge distribution. Therefore, while nuclear charge-radius measurements can help define the proton distribution, additional information (or assumptions) are required to define the neutron distribution.
While it is difficult to probe the neutron distribution directly, it is possible to measure the proton-neutron radius difference, $\Delta r_{np}=r_{rms}^{n}-r_{rms}^{p}$, via nuclear spectroscopy of antiprotonic atoms [32]. Presently, these data have large uncertainties, however, an approximately linear relationship between $\Delta r_{np}$ and $I\equiv(N-Z)/A$ is apparent. A global fit to the available data is given in reference [33]:
$\Delta r_{np}=[(0.82\pm 0.54)I-(0.02\pm 0.08)]\,\rm{fm}.$ (10)
Here we disregard the small uncertainty values given by Eq. (10), as we don’t believe the simple linear relationship captures the true scatter that exists in the data. However, we can use this formula to estimate that $\Delta r_{np}=0.050\,\rm{fm}$ for 70Ge ($I=0.09$) and $\Delta r_{np}=0.15\,\rm{fm}$ for 136Xe ($I=0.21$). All other nuclei in our study lie within this range of $I$ values. Therefore, following [24], we conservatively assume that the neutron rms radius lies within the range $r^{p}_{rms}<r^{n}_{rms}<r^{p}_{rms}+0.3$ fm.
@@
To define our empirical form factors we take $r^{p}_{rms}$ directly from electric-charge radii measurements, and use the range of $r^{n}_{rms}$ to define an uncertainty band. For the form factor we use the Klein-Nystrand (KN) [23] model, which has been previously used in CEvNS research (including by the COHERENT collaboration). The KN model is based on a hard sphere distribution with radius $r_{A}$ and a Yukawa potential of range $a_{k}$. The form factor is given by:
$F_{\text{KN}}(q)=3\frac{\sin(qr_{A})-qr_{A}\cos(qr_{A})}{(qr_{A})^{3}}\left[1+(qa_{k})^{2}\right]^{-1}.$ (11)
Here we take $a_{k}=0.7$ fm and obtain $r_{A}$ by equating the rms radius for the distribution with the proton rms charge radius:
$\langle r_{p}^{2}\rangle\equiv\langle r^{2}\rangle_{KN}=\frac{3}{5}r_{A}^{2}+6a_{k}^{2}.$ (12)
@@
The proton rms charge radii for the target nuclei are shown in table 1. The neutron rms radii are then taken to lie in the range $r_{\text{rms}}^{p}\leq r_{\text{rms}}^{n}\leq r_{\text{rms}}^{p}+0.3$ fm for Ge, I and Cs, whilst for Ar the upper limit is taken to be $r_{\text{rms}}^{p}+0.2$ fm.
II.3 Skyrme-Hartree-Fock distributions
@@
To model the weak-charge distribution of the nucleus more directly, we perform Skyrme-Hartree-Fock calculations to find the point-like nucleon distributions within the nucleus. The weak radius of the nucleus can then be calculated from:
$R_{W}^{2}=\frac{1}{Q_{\text{w}}}\left[ZQ_{\text{w}}^{p}(R_{p}^{2}+r_{\text{w},p}^{2})+(A-Z)Q_{\text{w}}^{n}(R_{n}^{2}+r_{\text{w},n}^{2})\right]$ (13)
where $r_{\text{w},p}=1.545(17)$ fm and $r_{\text{w},n}=0.8434(12)$ fm are the weak radii of the proton and neutron, respectively. In this equation we have dropped the flavor dependence for visual clarity. We take it as understood that all charges and radii are neutrino-flavor dependent and retain that dependence in all calculations.
@@
The nucleon densities were evaluated with the skyax code [34] assuming axial symmetry for several parametrizations of the Skyrme functional, namely SkM, SLy6, SV-min and UNEDF1 as defined in skyax. Pairing correlations are treated at the BCS level with a “density-dependent delta interaction” (DDDI) for SV-min and SkM, and with a “volume delta interaction” (VDI) for SLy6 and UNEDF1. The final theoretical values of $R_{p,n}$ are obtained from the average of the results for each interaction, while the standard deviations are used to evaluate the uncertainties. The full results for each interaction are compiled in the appendix (Table 6). From this we find the average and standard deviation for each isotope and use it to compute weak radii from Eq. (13). The results are summarised in Table 2.
@@
For simplicity we then parameterize the distribution via the symmetrized 2-parameter Fermi model [35], which has been shown to provide a reasonable fit to similar calculations [31]. The Fermi distribution is defined by:
@@
The (normalized) form factor is then:
@@
×$\times\frac{\pi qa}{\sinh(\pi qa)}\frac{1}{1+(\pi a/c)^{2}}.$
@@
Table 2: Skyrme-Hartree-Fock predictions of rms point-like proton and neutron radii and the resulting predicted rms Weak radius (for $\nu_{e}$).
@@
III Detailed nuclear treatment
@@
The full, bottom-up, calculation starts with the hadronic currents within the nucleus and using effective field theory, derives the Wilson coefficients for the nuclear couplings. This task has been performed for the CEvNS process, in full generality, in [18]. The resulting cross section is similar to Eq. (3), with an additional term for the axial form factor:
@@
However, the full nuclear form factor goes beyond the density distributions discussed previously. The Weak form factor becomes
@@
where $m_{N}$ is the nucleon mass and the remaining constants are provided in [18]. The weak form factor hence depends on the coherent and semi-coherent responses $\mathcal{F}^{M}_{N}$ and $\mathcal{F}^{\Phi^{\prime\prime}}_{N}$, respectively, where $N=\{p,n\}$ denotes protons and neutrons. The axial form factor is given by
@@
The correction to the leading spin-dependent (SD) coupling for the transverse SD response $\delta^{\prime}(\mathbf{q}^{2})$ is given in [18]. In this more-detailed treatment of the nucleus, the full weak radius receives additional contributions from spin-orbit effects and the strange electric-charge distribution:
@@
To make use of this formalism we need to calculate the nuclear responses $\mathcal{F}_{N}^{M}$, $\mathcal{F}_{N}^{\Phi^{\prime\prime}}$, and $\mathcal{F}_{N}^{\Sigma^{\prime}}$ for our chosen targets – for this we employ the nuclear shell model.
III.1 Evaluating the nuclear responses
Figure 1: Spin-independent weak form factors for CEvNS in the four targets relevant to current experiments. Bands show the error due to uncertainties in the neutron distribution (Charge) and from using different nuclear structure models (SHF and Shell model). The arrows indicate the COHERENT regions of interest (ROI).
The nuclear responses originate from a multipole decomposition of the lepton-nucleus amplitude, developed in [36, 37]. Here, they are defined as: